20191213 UT SX SOUL Commissioning on sky

SOUL SX

00:54 UT David points and collimates authorized for LUCI on left only

01:09 Bright Star, 8.32
  • bin1, 1700, 452
  • 8.32 catalog, 10.1 OCAM
  • 013805
  • 013849
  • 013931
  • 013400_jitter.jpg:
    013400_jitter.jpg
  • 013400_modal.jpg:
    013400_modal.jpg
  • 013400_gopt.jpg:
    013400_gopt.jpg

01:41 Dark 014151

01:45
  • 014508
  • 014623
  • 014703
  • SR = 0.34
  • 014623_psf.jpg:
    014623_psf.jpg

01:53 New AdSec Software to automatically turn on GOPT
  • Stop/Start AdSec Arbitrator
  • Did not start GOPT

02:42
  • bin 1, 1700, 362
  • seeing 0.9"
  • 8.5 catalog, 9.8 measured
  • 024249
  • 024352
  • 024507

02:46 Dark 024648

02:47
  • 024726
  • 024809
  • 024931
  • 025057 (25)
  • SR = 0.59
  • 024931_jitter.jpg:
    024931_jitter.jpg
  • 024931_gopt.jpg:
    024931_gopt.jpg
  • 024931_modal.jpg:
    024931_modal.jpg
  • 024931_psf.jpg:
    024931_psf.jpg

02:52 NCPA On
  • 025226
  • 025319
  • 025414
  • 025625
  • 025316_psf.jpg:
    025316_psf.jpg

03:56 16.4 mag
  • bad seeing up to 3"
  • 035612 (No LUCI)
  • 035756 (with full frame LUCI 34)
  • 040357
  • 040511
  • 040612

04:07 Dark 040839

04:16
  • 041638
  • 041747

04:20 Dark 042018

04:23
  • 042338
  • 042611

04:49 New amplitude on GOPT signal
  • GOPT = 0.07, sm

04:51 msgdm and almost all processes go down
  • loose all GUI's
  • w_stop, w_start
  • click Chill on OCAM

05:06
  • GOPT 0.03
  • 050621 2.51 sec
  • 050727
05:08 WFS processes die * while bringing them up, three AdSec processes died
[AOeng@sxadsec 13]$ adsc_check

msgdrtdb             running
adamhousekeeper      NOT running
mirrorctrl           running but NOT connected to Msgd
pinger               running but NOT connected to Msgd
masterdiagnostic     running but NOT connected to Msgd
idlctrl              NOT running
housekeeper          running but NOT connected to Msgd
fastdiagn            NOT running
adsecarb             running but NOT connected to Msgd
AOARB                running but NOT connected to Msgd
varsmonitor          running but NOT connected to Msgd
anemometermon        running but NOT connected to Msgd

05:10 adsc_stop/start

05:20 WFS Power status show all power off
  • New HW GUI show all power on but RR stuff
  • power_status_wrong.jpg:
    power_status_wrong.jpg

05:30 Get "AcquireRef{FLAO}[825] - Command interrupted" error on RunAO
  • Resend RunAO and get same error
  • Send Preset and RunAO, success

05:35 Center Pupil fail
  • Rerotator was not on. Should have been turned on by PresetAO
  • Why not?

05:47
  • bin 1, 516, 600
  • catalog 12.5, see 13.1
  • seeing 0.8"
  • 054748
  • 054845
  • 055044
  • 055044_jitter.jpg:
    055044_jitter.jpg
  • 055044_modal.jpg:
    055044_modal.jpg
  • 055044_gopt.jpg:
    055044_gopt.jpg

05:52 Dark (53) 055337

05:54
  • 055436
  • 055620
  • 055735
  • 060117 move 0.5" away from bad pixel
  • 060218 (59)
  • 060339

06:06 NCPA On
  • 060633
  • 060742
  • 060839

06:41 13.3 catalog, 13.2 measured
  • bin 1, 950, 600
  • seeing 1.1"
  • 064148 (no luci, change number)
  • 064311 (67)
  • 064525
  • 064651
  • 064651_jitter.jpg:
    064651_jitter.jpg
  • 064651_modal.jpg:
    064651_modal.jpg
  • 064651_gopt.jpg:
    064651_gopt.jpg

06:48 Dark 064924

06:50
  • 065001
  • 065053
  • 065143

06:52 NCPA On
  • 065253 (75)
  • 065344
  • 065607
  • 065702

OSCO

07:26 OSCO Script
  • 7.46 catalog, 8.4 mag measured
  • Bin 1, 1700, 100
  • seeing 0.7"-1.0"
  • 073745
  • 073838
  • 073906
  • luci 79-98
  • Looked like AO loop overdriving at 230 Hz and a bit at 127 Hz

07:41 IDL Preset to AO173
  • 7.5 catalog,
  • bin 1, 1700, 100
  • seeing 1.0"
  • 075104
  • 075121
  • luci 99-109
  • Looked like AO loop overdriving at 230 Hz and a bit at 127 Hz
  • At end of OSCO script the AO loop is left in Pause. Should Resume
Fri Dec 13 07:56:14.595 2019 left Offset mount delay for LUCIFER: 7230 milliseconds (minimum 0)
Fri Dec 13 07:56:14.756 2019 left Offset alert for LUCIFER: Pause{FLAO}[820]
Fri Dec 13 07:56:14.769 2019 left Offset alert for LUCIFER: WFS loop is not paused
Fri Dec 13 07:56:14.786 2019 left Offset alert for LUCIFER: OffsetXY{FLAO}[971]
Fri Dec 13 07:56:14.803 2019 left Offset command complete with warnings for LUCIFER, Sound=offsetWarning.wav
  • 073745_modal.jpg:
    073745_modal.jpg
  • 073745_jitter_1.jpg:
    073745_jitter_1.jpg
  • 073745_jitter_2.jpg:
    073745_jitter_2.jpg

-- DougMiller - 13 Dec 2019

  • Time history of PSD peak at 220 Hz:
    freq.mode2.20191213telescope.jpg

  • OVMS TT Prediction Time Series and PSD:
    ovms.mode2.20191213telescope.jpg
I Attachment Action Size Date Who Comment
013400_gopt.jpgjpg 013400_gopt.jpg manage 28 K 13 Dec 2019 - 01:37 DougMiller  
013400_jitter.jpgjpg 013400_jitter.jpg manage 30 K 13 Dec 2019 - 01:37 DougMiller  
013400_modal.jpgjpg 013400_modal.jpg manage 32 K 13 Dec 2019 - 01:37 DougMiller  
014623_psf.jpgjpg 014623_psf.jpg manage 36 K 13 Dec 2019 - 02:24 DougMiller  
024931_gopt.jpgjpg 024931_gopt.jpg manage 29 K 13 Dec 2019 - 02:55 DougMiller  
024931_jitter.jpgjpg 024931_jitter.jpg manage 31 K 13 Dec 2019 - 02:54 DougMiller  
024931_modal.jpgjpg 024931_modal.jpg manage 32 K 13 Dec 2019 - 02:55 DougMiller  
024931_psf.jpgjpg 024931_psf.jpg manage 38 K 13 Dec 2019 - 03:02 DougMiller  
025316_psf.jpgjpg 025316_psf.jpg manage 38 K 13 Dec 2019 - 03:02 DougMiller  
055044_gopt.jpgjpg 055044_gopt.jpg manage 28 K 13 Dec 2019 - 07:07 DougMiller  
055044_jitter.jpgjpg 055044_jitter.jpg manage 27 K 13 Dec 2019 - 07:07 DougMiller  
055044_modal.jpgjpg 055044_modal.jpg manage 29 K 13 Dec 2019 - 07:07 DougMiller  
064651_gopt.jpgjpg 064651_gopt.jpg manage 28 K 13 Dec 2019 - 07:06 DougMiller  
064651_jitter.jpgjpg 064651_jitter.jpg manage 26 K 13 Dec 2019 - 07:06 DougMiller  
064651_modal.jpgjpg 064651_modal.jpg manage 30 K 13 Dec 2019 - 07:06 DougMiller  
073745_jitter_1.jpgjpg 073745_jitter_1.jpg manage 26 K 13 Dec 2019 - 08:01 DougMiller  
073745_jitter_2.jpgjpg 073745_jitter_2.jpg manage 24 K 13 Dec 2019 - 08:02 DougMiller  
073745_modal.jpgjpg 073745_modal.jpg manage 34 K 13 Dec 2019 - 08:01 DougMiller  
freq.mode2.20191213telescope.jpgjpg freq.mode2.20191213telescope.jpg manage 535 K 13 Dec 2019 - 21:55 JohnHill Time history of PSD peak at 220 Hz
ovms.mode2.20191213telescope.jpgjpg ovms.mode2.20191213telescope.jpg manage 1 MB 13 Dec 2019 - 21:56 JohnHill OVMS TT Prediction Time Series and PSD
power_status_wrong.jpgjpg power_status_wrong.jpg manage 115 K 13 Dec 2019 - 05:24 DougMiller  
Topic revision: r5 - 13 Dec 2019, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback